# Diffusion Transformer Captures Spatial-Temporal Dependencies: A Theory for Gaussian Process Data

### v1

@@ B.3.3 Formal Proof of Lemma 6
$\mathbf{s}(\mathbf{v}_{t})=(\alpha_{t}^{2}(\bm{\Gamma}\otimes\bm{\Sigma})+\sigma_{t}^{2}\mathbf{I})^{-1}(\mathbf{v}_{t}-\alpha_{t}\bm{\mu}).$
@@
Size of Transformer Architecture
We summarize our construction by characterizing the size of the architecture of the entire network:
  • The input to the transformer is of dimension $D\times N$ with $D=9d+d_{e}+d_{t}+1$.
  • The number of transformer blocks
    $L=L_{\mu}+(2+L_{\rm mult})K=\mathcal{O}(\kappa_{t_{0}}\log(R_{0}Nd/\epsilon)^{2}).$
  • The number of attention heads in each block is bounded by
    $M=4J=\mathcal{O}{\left({\left(\ell\log(N\ell\left\lVert\bm{\Sigma}\right\rVert_{\rm F}/\sigma_{t_{0}})\right)}^{1/v}\right)}.$
  • By the construction above and Lemma 9, 10 and our upper bound on the input in each transformer block, we conclude that the Frobenius norm of each parameter is bounded by
    $B=\mathcal{O}{\left(\log(R_{0}Nd/(\epsilon\sigma_{t_{0}}))\sqrt{N}R_{0}\sigma^{-2}_{t_{0}}(r^{2}+\left\lVert\bm{\Sigma}\right\rVert_{\infty})\right)}.$
  • The output range is $R_{t}=2R_{0}\sigma_{t}^{-1}$.
The proof is complete.
@@ B.4 Omitted Proofs of Lemmas about the Score Function
Suppose $\mathbf{v}_{t}\sim P_{t}=\mathcal{N}(\alpha_{t}\bm{\mu},\alpha_{t}^{2}({\bm{\Gamma}}\otimes\bm{\Sigma})+\sigma_{t}^{2}\mathbf{I})$. Denote $\bm{\Phi}_{t}=\alpha_{t}^{2}({\bm{\Gamma}}\otimes\bm{\Sigma})+\sigma_{t}^{2}\mathbf{I}$ and $\bar{\bm{\Phi}}_{t}=\alpha_{t}^{2}(\bar{\bm{\Gamma}}\otimes\bm{\Sigma})+\sigma_{t}^{2}\mathbf{I}$. Then we have ${\bm{\Gamma}}_{t}=\bar{\bm{\Phi}}_{t}+\alpha_{t}^{2}(\Delta\bm{\Gamma}\otimes\bm{\Sigma})$. Moreover, by Lemma 4, choosing $J=\mathcal{O}((\log(N\left\lVert\bm{\Sigma}\right\rVert_{\rm F}/(\epsilon\sigma_{t})))$ ensures that $\left\lVert\Delta\bm{\Gamma}\right\rVert_{\rm F}\leq\epsilon\left\lVert\bm{\Sigma}\right\rVert^{-2}_{\rm F}\sigma_{t}^{2}$. Thus, we have
@@
Note that if $\mathbf{v}_{t}\sim P_{t}$ we can write $\mathbf{v}_{t}=\alpha_{t}\bm{\mu}+\bm{\Phi}_{t}^{1/2}\mathbf{z}$, where $\bm{\Phi}_{t}=\alpha_{t}^{2}\bm{\Gamma}\otimes\bm{\Sigma}+\sigma_{t}^{2}\mathbf{I}$ and $\mathbf{z}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$. Thus, the score function can be written as
$\nabla\log p_{t}(\mathbf{v}_{t})=\bm{\Phi}_{t}^{-1}(\mathbf{v}_{t}-\alpha_{t}\bm{\mu})=\bm{\Phi}_{t}^{-1/2}\mathbf{z}.$
@@ B.5 Results on Transformers with Softmax Activation
$f_{{\rm GD,softmax}}{\left(\mathbf{y}_{i}\right)}$ $=\Bigg{[}\mathbf{x}_{i}^{\top},\bm{e}_{i}^{\top},\bm{\phi}^{\top}(t),\quad\mathbf{s}_{i}^{\top}-\eta{\left(\sum_{j=1}^{N}(\alpha_{t}^{2}{\bm{\Gamma}}_{ij}\bm{\Sigma}+\sigma_{t}^{2}\mathbf{I})\mathbf{s}_{j}^{\top}-(\mathbf{x}_{i}-\alpha_{t}\bm{\mu}_{i})^{\top}\right)}+\bm{\epsilon}_{i},$
@@
$\mathbf{s}_{i}^{(k+1)}$ $=\mathbf{s}_{i}^{(k)}-\eta_{t}\left[\sum_{j=1}^{N}(\alpha_{t}^{2}{\bm{\Gamma}}_{ij}\bm{\Sigma})\mathbf{s}_{j}^{(k)}+\sigma^{2}_{t}\mathbf{s}_{i}^{(k)}-(\mathbf{x}_{i,t}-\alpha_{t}\bm{\mu}_{i})\right]$
$=\mathbf{s}_{i}^{(k)}-\underbrace{\eta_{t}\sum_{j=1}^{N}\alpha_{t}^{2}{\bm{\Gamma}}_{ij}\bm{\Sigma}\mathbf{s}_{j}^{(k)}}_{(A)}+\underbrace{\left(-\eta_{t}\sigma_{t}^{2}\mathbf{s}_{i}^{(k)}+\eta_{t}(\mathbf{x}_{i,t}-\alpha_{t}\bm{\mu}_{i})\right)}_{(B)}.$ (14)
@@
$(A)$ $=\eta\alpha^{2}\sum_{j=1}^{N}\exp(-\left\lVert\mathbf{e}_{i}-\mathbf{e}_{j}\right\rVert_{2}^{2}/\ell)\bm{\Sigma}\mathbf{s}_{j}$
@@
Size of Transformer Architecture
We summarize our construction by characterizing the size of the architecture of $f_{\rm GD,softmax}$:
  • The input to the transformer is of dimension $D\times N$ with $D=9d+d_{e}+d_{t}+1$.
  • The number of transformer blocks
    $L=2+L_{\rm mult}=\mathcal{O}(\log(\left\lVert\mathbf{x}\right\rVert_{\infty}\left\lVert\mathbf{s}\right\rVert_{\infty}Nd/\epsilon)^{2}).$
  • Only one attention head is needed, i.e., $M=1$.
  • By the construction above, we conclude that the norm of each parameter is bounded by $\mathcal{O}(N^{2}+d+\left\lVert\bm{\Sigma}\right\rVert_{F}+r^{2}))$.
The proof is complete. ∎ With Lemma 11 and following the proof in Appendix B while omitting the part of truncating $\bm{\Gamma}$, i.e., setting $J=N$, we could similarly construct a softmax-transformers as
@@ C.1 Proof of Theorem 2
$\widehat{\mathbf{s}}_{t}=\arg\min_{s_{t}\in\mathcal{F}}\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}_{\mathbf{v}_{t}\mid\mathbf{v}_{0}^{(i)}}\left\|\mathbf{s}_{t}(\mathbf{v}_{t})-\nabla\log p_{t}(\mathbf{v}_{t}\mid\mathbf{v}_{0}^{(i)})\right\|_{2}^{2}.$
@@
$\int_{\mathbf{x}}p_{t}(\mathbf{x})\left\lVert(\mathbf{s}-\mathbf{s}^{\prime})(\mathbf{x},t)\right\rVert_{2}^{2}{\rm d}\mathbf{v}_{t}\leq C$ (16)
@@
$\text{KL}\left(p_{T}\|p^{\prime}_{T}\right)\leq\int_{0}^{T}\frac{1}{2}\int_{\mathbf{x}}p_{t}(\mathbf{x})\left\lVert(\mathbf{s}-\mathbf{s}^{\prime})(\mathbf{x},t)\right\rVert_{2}^{2}{\rm d}\mathbf{x}{\rm d}t.$
@@
$\int_{\mathbf{v}_{t}}p_{t}(\mathbf{v}_{t})\left\lVert\mathbf{s}_{t}(\mathbf{v}_{t})-\nabla\log p_{t}(\mathbf{v}_{t})\right\rVert_{2}^{2}{\rm d}\mathbf{v}_{t}$ $\lesssim\int_{\mathbf{v}_{t}}p_{t}(\mathbf{v}_{t}){\left(\frac{\left\lVert\mathbf{v}_{t}-\alpha_{t}\bm{\mu}\right\rVert_{2}^{2}}{\sigma^{4}_{t}}+\frac{C}{\sigma_{t}^{2}}\right)}{\rm d}\mathbf{v}_{t}\lesssim\frac{1}{\sigma^{2}_{t}}.$
@@
$\lesssim\exp(-T)+\sqrt{\int_{t_{0}}^{T}\frac{1}{2}\int_{\mathbf{v}_{t}}p_{t}(\mathbf{v}_{t})\left\lVert\widehat{\mathbf{s}}(\mathbf{v}_{t},\mathbf{y},t)-\nabla\log p_{t}(\mathbf{v}_{t})\right\rVert_{2}^{2}{\rm d}\mathbf{v}_{t}{\rm d}t}.$ (17)
@@ C.2.1 Additional Notations
$\ell(\widehat{\mathbf{s}}_{t}):=\int_{t_{0}}^{T}{\rm d}t\cdot\mathbb{E}_{\mathbf{v}_{t}\sim p_{t}}\left\|\widehat{\mathbf{s}}_{t}(\mathbf{v}_{t})-\nabla\log p_{t}(\mathbf{v}_{t})\right\|_{2}^{2}$
@@
^$\widehat{\ell}(\widehat{\mathbf{s}}_{t})$ $=\frac{1}{n}\sum_{i=1}^{n}\int_{t_{0}}^{T}{\rm d}t\cdot\mathbb{E}_{\mathbf{v}_{t}\mid\mathbf{v}_{0}^{(i)}}\left\|\widehat{\mathbf{s}}_{t}(\mathbf{v}_{t})+\frac{\mathbf{v}_{t}-\alpha_{t}\mathbf{v}_{0}^{(i)}}{\sigma_{t}^{2}}\right\|_{2}^{2}.$
@@
$\widehat{\mathbf{s}}_{t}=\arg\min_{s_{t}\in\mathcal{F}}\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}_{\mathbf{v}_{t}\mid\mathbf{v}_{0}^{(i)}}\left\|s_{t}(\mathbf{v}_{t})-\nabla\log p_{t}(\mathbf{v}_{t}\mid\mathbf{v}_{0}^{(i)})\right\|_{2}^{2}.$
@@ C.2.2 Proof of Proposition 2
$h(\widehat{\mathbf{s}},\mathbf{v}_{0}):=\int_{t_{0}}^{T}{\rm d}t\cdot\mathbb{E}_{\mathbf{v}_{t}\mid\mathbf{v}_{0}}\left\|\widehat{\mathbf{s}}_{t}(\mathbf{v}_{t})+\frac{\mathbf{v}_{t}-\alpha_{t}\mathbf{v}_{0}}{\sigma_{t}^{2}}\right\|_{2}^{2}\leq 2Nd(T+\log(1/t_{0}))(R_{s}^{2}+1).$ (18)
@@
^$\widehat{\ell}^{\rm trunc}(\widehat{\mathbf{s}}_{t})$ $:=\frac{1}{n}\sum_{i=1}^{n}\int_{t_{0}}^{T}{\rm d}t\cdot\mathbb{E}_{\mathbf{v}_{t}\mid\mathbf{v}_{0}^{(i)}}{\left[\left\|\widehat{\mathbf{s}}_{t}(\mathbf{v}_{t})+\frac{\mathbf{v}_{t}-\alpha_{t}\mathbf{v}_{0}^{(i)}}{\sigma_{t}^{2}}\right\|_{2}^{2}{\mathbf{1}\left\{\left\lVert v_{0}^{(i)}\right\rVert_{2}\leq R_{0}\right\}}\right]}.$
@@
${\left|\mathbb{E}_{\mathbf{v}_{0}}{\left[\int_{t_{0}}^{T}{\rm d}t\cdot\mathbb{E}_{\mathbf{v}_{t}\mid\mathbf{v}_{0}^{(i)}}\left\|\widehat{\mathbf{s}}_{t}(\mathbf{v}_{t})+\frac{\mathbf{v}_{t}-\alpha_{t}\mathbf{v}_{0}}{\sigma_{t}^{2}}\right\|_{2}^{2}{\mathbf{1}\left\{\left\lVert\mathbf{v}_{0}\right\rVert\geq R_{0}\right\}}\right]}\right|}\lesssim(T+\log(1/t_{0}))\epsilon.$

### v2

@@ B.3.3 Formal Proof of Lemma 6
$\mathbf{s}(\mathbf{v}_{t})=-(\alpha_{t}^{2}(\bm{\Gamma}\otimes\bm{\Sigma})+\sigma_{t}^{2}\mathbf{I})^{-1}(\mathbf{v}_{t}-\alpha_{t}\bm{\mu}).$
@@
The Overall Size of Transformer Architecture for Approximating The Score Function
We combine the network sizes in Table 1 and Table 2, which gives rise to a characterization of the overall network size for approximating the score function of Gaussian process data. The result is summarized in the following table.
Table 3: Overall transformer network size for approximating the score function
Input dimension $D\times N$ with $D=9d+d_{e}+d_{t}+1$
# of blocks $L$ $L_{\mu}+(2+L_{\rm mult})K=\mathcal{O}(\kappa_{t_{0}}\log(R_{0}Nd/\epsilon)^{2})$
# of attention heads $M$ $4J=\mathcal{O}{\left({\left(\ell\log(N\ell\left\lVert\bm{\Sigma}\right\rVert_{\rm F}/\sigma_{t_{0}})\right)}^{1/v}\right)}$
Parameter bound $B$ $\mathcal{O}{\left(\log(R_{0}Nd/(\epsilon\sigma_{t_{0}}))\sqrt{N}R_{0}\sigma^{-2}_{t_{0}}(r^{2}+\left\lVert\bm{\Sigma}\right\rVert_{\infty})\right)}$
Output range $R_{t}$ $2R_{0}\sigma_{t}^{-1}$
Note that to obtain the bound on $L$ and $M$, we substitute the choice of $J$ and $K$ into the network sizes in Tables 1 and 2. The parameter bound is obtained by explicitly evaluating the magnitude $\left\lVert\mathbf{s}\right\rVert_{\infty}$ of the score function in Table 2. The proof is complete.
@@ B.4 Omitted Proofs of Lemmas about the Score Function
Suppose $\mathbf{v}_{t}\sim P_{t}=\mathcal{N}(\alpha_{t}\bm{\mu},\alpha_{t}^{2}({\bm{\Gamma}}\otimes\bm{\Sigma})+\sigma_{t}^{2}\mathbf{I}))$. Denote $\bm{\Phi}_{t}=\alpha_{t}^{2}({\bm{\Gamma}}\otimes\bm{\Sigma})+\sigma_{t}^{2}\mathbf{I}$ and $\bar{\bm{\Phi}}_{t}=\alpha_{t}^{2}(\bar{\bm{\Gamma}}\otimes\bm{\Sigma})+\sigma_{t}^{2}\mathbf{I}$. Then we have ${\bm{\Gamma}}_{t}=\bar{\bm{\Phi}}_{t}+\alpha_{t}^{2}(\Delta\bm{\Gamma}\otimes\bm{\Sigma})$. Moreover, by Lemma 4, choosing $J=\mathcal{O}((\log(N\left\lVert\bm{\Sigma}\right\rVert_{\rm F}/(\epsilon\sigma_{t})))$ ensures that $\left\lVert\Delta\bm{\Gamma}\right\rVert_{\rm F}\leq\epsilon\left\lVert\bm{\Sigma}\right\rVert^{-2}_{\rm F}\sigma_{t}^{2}$. Thus, we have
@@
Note that if $\mathbf{v}_{t}\sim P_{t}$ we can write $\mathbf{v}_{t}=\alpha_{t}\bm{\mu}-\bm{\Phi}_{t}^{1/2}\mathbf{z}$, where $\bm{\Phi}_{t}=\alpha_{t}^{2}\bm{\Gamma}\otimes\bm{\Sigma}+\sigma_{t}^{2}\mathbf{I}$ and $\mathbf{z}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$. Thus, the score function can be written as
$\nabla\log p_{t}(\mathbf{v}_{t})=-\bm{\Phi}_{t}^{-1}(\mathbf{v}_{t}-\alpha_{t}\bm{\mu})=\bm{\Phi}_{t}^{-1/2}\mathbf{z}.$
@@ B.5 Results on Transformers with Softmax Activation
$f_{{\rm GD,softmax}}{\left(\mathbf{y}_{i}\right)}$ $=\Bigg{[}\mathbf{x}_{i}^{\top},\bm{e}_{i}^{\top},\bm{\phi}^{\top}(t),\quad\mathbf{s}_{i}^{\top}-\eta_{t}{\left(\sum_{j=1}^{N}(\alpha_{t}^{2}{\bm{\Gamma}}_{ij}\bm{\Sigma}+\sigma_{t}^{2}\mathbf{I})\mathbf{s}_{j}^{\top}-(\mathbf{x}_{i}-\alpha_{t}\bm{\mu}_{i})^{\top}\right)}+\bm{\epsilon}_{i},$
@@
$\mathbf{s}_{i}^{(k+1)}$ $=\mathbf{s}_{i}^{(k)}-\eta_{t}\left[\sum_{j=1}^{N}(\alpha_{t}^{2}{\bm{\Gamma}}_{ij}\bm{\Sigma})\mathbf{s}_{j}^{(k)}+\sigma^{2}_{t}\mathbf{s}_{i}^{(k)}+(\mathbf{x}_{i,t}-\alpha_{t}\bm{\mu}_{i})\right]$
$=\mathbf{s}_{i}^{(k)}-\underbrace{\eta_{t}\sum_{j=1}^{N}\alpha_{t}^{2}{\bm{\Gamma}}_{ij}\bm{\Sigma}\mathbf{s}_{j}^{(k)}}_{(A)}-\underbrace{\left(\eta_{t}\sigma_{t}^{2}\mathbf{s}_{i}^{(k)}+\eta_{t}(\mathbf{x}_{i,t}-\alpha_{t}\bm{\mu}_{i})\right)}_{(B)}.$ (14)
@@
$(A)$ $=\eta\alpha^{2}\sum_{j=1}^{N}\exp(-\left\lVert\mathbf{e}_{i}-\mathbf{e}_{j}\right\rVert^{2}/\ell)\bm{\Sigma}\mathbf{s}_{j}$
@@
Size of Transformer Architecture with Softmax Activation
We list here the size of transformer architecture equipped with the softmax activation. Readers may draw a quick comparison with the size of ReLU transformer in Table 3.
Table 4: Size of transformer with Softmax activation for approximating score function (with Gaussian covariance function)
Input dimension $D\times N$ with $D=9d+d_{e}+d_{t}+1$
# of blocks $L$ $2+L_{\rm mult}$
# of attention heads $M$ $1$
Parameter bound $B$ $\mathcal{O}(N^{2}+d+\left\lVert\bm{\Sigma}\right\rVert_{\rm F}+r^{2}))$
Output range $R_{t}$ $2R_{0}\sigma_{t}^{-1}$
It is observed that using Softmax activation leads to a reduced number of transformer blocks. The reason behind is that the Gaussian covariance function can be approximately represented by one attention block. The proof is complete.
∎ With Lemma 11 and following the proof in Appendix B while omitting the part of truncating $\bm{\Gamma}$, i.e., setting $J=N$, we could similarly construct a softmax-transformers as
@@ C.1 Proof of Theorem 2
$\widehat{\mathbf{s}}_{t}=\arg\min_{s_{t}\in\mathcal{F}}\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}_{\mathbf{v}_{t}\mid\mathbf{v}_{0}^{(i)}}\left\|\mathbf{s}_{t}(\mathbf{v}_{t})-\nabla\log p_{t}(\mathbf{v}_{t}\mid\mathbf{v}_{0}^{(i)})\right\|^{2}_{2}.$
@@
$\int_{\mathbf{x}}p_{t}(\mathbf{x})\left\lVert(\mathbf{s}-\mathbf{s}^{\prime})(\mathbf{x},t)\right\rVert^{2}_{2}{\rm d}\mathbf{v}_{t}\leq C$ (16)
@@
$\text{KL}\left(p_{T}\|p^{\prime}_{T}\right)\leq\int_{0}^{T}\frac{1}{2}\int_{\mathbf{x}}p_{t}(\mathbf{x})\left\lVert(\mathbf{s}-\mathbf{s}^{\prime})(\mathbf{x},t)\right\rVert^{2}_{2}{\rm d}\mathbf{x}{\rm d}t.$
@@
$\int_{\mathbf{v}_{t}}p_{t}(\mathbf{v}_{t})\left\lVert\mathbf{s}_{t}(\mathbf{v}_{t})-\nabla\log p_{t}(\mathbf{v}_{t})\right\rVert^{2}_{2}{\rm d}\mathbf{v}_{t}$ $\lesssim\int_{\mathbf{v}_{t}}p_{t}(\mathbf{v}_{t}){\left(\frac{\left\lVert\mathbf{v}_{t}-\alpha_{t}\bm{\mu}\right\rVert^{2}_{2}}{\sigma^{4}_{t}}+\frac{C}{\sigma_{t}^{2}}\right)}{\rm d}\mathbf{v}_{t}\lesssim\frac{1}{\sigma^{2}_{t}}.$
@@
$\lesssim\exp(-T)+\sqrt{\int_{t_{0}}^{T}\frac{1}{2}\int_{\mathbf{v}_{t}}p_{t}(\mathbf{v}_{t})\left\lVert\widehat{\mathbf{s}}(\mathbf{v}_{t},\mathbf{y},t)-\nabla\log p_{t}(\mathbf{v}_{t})\right\rVert^{2}_{2}{\rm d}\mathbf{v}_{t}{\rm d}t}.$ (17)
@@ C.2.1 Additional Notations
$\ell(\widehat{\mathbf{s}}_{t}):=\int_{t_{0}}^{T}{\rm d}t\cdot\mathbb{E}_{\mathbf{v}_{t}\sim p_{t}}\left\|\widehat{\mathbf{s}}_{t}(\mathbf{v}_{t})-\nabla\log p_{t}(\mathbf{v}_{t})\right\|^{2}_{2}$
@@
^$\widehat{\ell}(\widehat{\mathbf{s}}_{t})$ $=\frac{1}{n}\sum_{i=1}^{n}\int_{t_{0}}^{T}{\rm d}t\cdot\mathbb{E}_{\mathbf{v}_{t}\mid\mathbf{v}_{0}^{(i)}}\left\|\widehat{\mathbf{s}}_{t}(\mathbf{v}_{t})+\frac{\mathbf{v}_{t}-\alpha_{t}\mathbf{v}_{0}^{(i)}}{\sigma_{t}^{2}}\right\|^{2}_{2}.$
@@
$\widehat{\mathbf{s}}_{t}=\arg\min_{s_{t}\in\mathcal{F}}\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}_{\mathbf{v}_{t}\mid\mathbf{v}_{0}^{(i)}}\left\|s_{t}(\mathbf{v}_{t})-\nabla\log p_{t}(\mathbf{v}_{t}\mid\mathbf{v}_{0}^{(i)})\right\|^{2}_{2}.$
@@ C.2.2 Proof of Proposition 2
$h(\widehat{\mathbf{s}},\mathbf{v}_{0}):=\int_{t_{0}}^{T}{\rm d}t\cdot\mathbb{E}_{\mathbf{v}_{t}\mid\mathbf{v}_{0}}\left\|\widehat{\mathbf{s}}_{t}(\mathbf{v}_{t})+\frac{\mathbf{v}_{t}-\alpha_{t}\mathbf{v}_{0}}{\sigma_{t}^{2}}\right\|^{2}_{2}\leq 2Nd(T+\log(1/t_{0}))(R_{s}^{2}+1).$ (18)
@@
^$\widehat{\ell}^{\rm trunc}(\widehat{\mathbf{s}}_{t})$ $:=\frac{1}{n}\sum_{i=1}^{n}\int_{t_{0}}^{T}{\rm d}t\cdot\mathbb{E}_{\mathbf{v}_{t}\mid\mathbf{v}_{0}^{(i)}}{\left[\left\|\widehat{\mathbf{s}}_{t}(\mathbf{v}_{t})+\frac{\mathbf{v}_{t}-\alpha_{t}\mathbf{v}_{0}^{(i)}}{\sigma_{t}^{2}}\right\|^{2}_{2}{\mathbf{1}\left\{\left\lVert v_{0}^{(i)}\right\rVert_{2}\leq R_{0}\right\}}\right]}.$
@@
${\left|\mathbb{E}_{\mathbf{v}_{0}}{\left[\int_{t_{0}}^{T}{\rm d}t\cdot\mathbb{E}_{\mathbf{v}_{t}\mid\mathbf{v}_{0}^{(i)}}\left\|\widehat{\mathbf{s}}_{t}(\mathbf{v}_{t})+\frac{\mathbf{v}_{t}-\alpha_{t}\mathbf{v}_{0}}{\sigma_{t}^{2}}\right\|^{2}_{2}{\mathbf{1}\left\{\left\lVert\mathbf{v}_{0}\right\rVert\geq R_{0}\right\}}\right]}\right|}\lesssim(T+\log(1/t_{0}))\epsilon.$